if 90 is added to a number the result is 36 less than 3 times that number.What is the number
Answer by umail08(87)

Let x = the number:
x + 90 … 90 is added to a number
3x … 3 times that number
3x -36 … 36 less than 3 times that number
x + 90 = 3x – 36 …this is the equation
x – 3x + 90 = 3x – 3x – 36
-2x + 90 = -36
-2x + 90 – 90 = -36 – 90
-2x = -126
-2x/-2 = -126/-2
x = 63